Paradise Corporation budgets on an annual basis for its fiscal year. The following beginning and ending inventory levels (in units) are planned for next year.

  Beginning Inventory Ending Inventory
  Raw material* 56,000 66,000
  Finished goods 96,000 66,000

*Three pounds of raw material are needed to produce each unit of finished product.

 

If Paradise Corporation plans to sell 560,000 units during next year, the number of units it would have to manufacture during the year would be:

504,000 units

560,000 units

590,000 units

530,000 units

2-

Morie Corporation is working on its direct labor budget for the next two months. Each unit of output requires 0.75 direct labor-hours. The direct labor rate is $10.00 per direct labor-hour. The production budget calls for producing 7,100 units in March and 6,900 units in April. The company guarantees its direct labor workers a 40-hour paid work week. With the number of workers currently employed, that means that the company is committed to paying its direct labor work force for at least 5,480 hours in total each month even if there is not enough work to keep them busy. What would be the total combined direct labor cost for the two months?

$109,600.00

$120,750.00

$105,000.00

$106,550.00

3-

Laurey Inc. is working on its cash budget for May. The budgeted beginning cash balance is $51,000. Budgeted cash receipts total $135,000 and budgeted cash disbursements total $130,000. The desired ending cash balance is $72,000. To attain its desired ending cash balance for May, the company needs to borrow:

$128,000

$0

$72,000

$16,000

4-

Richards Corporation has the following budgeted sales for the first half of next year:

  

  Cash Sales Credit Sales
  January $80,000        $180,000     
  February $85,000        $200,000     
  March $36,000        $160,000     
  April $41,000        $136,000     
  May $51,000        $230,000     
  June $110,000        $260,000     

  

The company is in the process of preparing a cash budget and must determine the expected cash collections by month. To this end, the following information has been assembled:

  

  Collections on credit sales:
  55% in month of sales
  30% in month of following sales
  15.0% in second month following sales

  

The accounts receivable balance on January 1 is $68,000. Of this amount, $54,000 represents uncollected December sales and $14,000 represents uncollected November sales.

The total cash collected during January would be:

$300,000

$241,000

$120,000

$229,000

5-

Richards Corporation has the following budgeted sales for the first half of next year:

  

  Cash Sales Credit Sales
  January $70,000        $170,000     
  February $75,000        $190,000     
  March $54,000        $150,000     
  April $59,000        $154,000     
  May $69,000        $220,000     
  June $100,000        $440,000     

  

The company is in the process of preparing a cash budget and must determine the expected cash collections by month. To this end, the following information has been assembled:

  

  Collections on credit sales:
  40% in month of sales
  30% in month of following sales
  30% in second month following sales
The accounts receivable balance on January 1 is $84,000. Of this amount, $48,000 represents uncollected December sales and $36,000 represents uncollected November sales.

What is the budgeted accounts receivable balance on May 30?

$178,200

$48,000

$226,200

$132,000
Answer #1
Concepts and reason

The concepts used to solve the given problem are financial statements and budgeting.

Financial Statement: The statements that are prepared for a specific period of time, comprising the financial information of the organization are known as financial statements. It includes statement of income which shows the profitability of the business, balance sheet which shows the monetary position with regards to assets, liabilities and capital, and cash flow statement which represents flow of cash for the accounting period.

Budgeting: A plan which is created by a company in order to determine how its resources are to be used to achieve the targeted profits and growth of the company is called budget. The process of creating such a plan is known as budgeting.

Fundamentals

Inventory: It refers to the stock of goods held by an entity for carrying on business. Inventory constitutes raw material stored in storeroom, work in progress used in production and finished goods available for sale.

Finished Goods (FG): The raw materials are processed and such goods which completes the process of manufacturing, but are not yet sold to the customer, are known as finished goods. It also includes goods in a complete form have been purchased for the purposes of sale.

Raw Material (RM): The material of a company which is unprocessed and will be used to manufacture finished products are called raw materials. They may be direct or indirect.

Direct Labor: The manpower which is employed to produce or manufacture the goods is called direct labor. They do not perform functions like administration, selling, maintenance and like.

Direct Labor-Hours: Direct labor hours is defined as hours spent by the labor to produce or manufacture the goods and is considered as one of the basis of allocation of overheads.

Cash: Cash is a legal tender which is a current asset of the company. It is the utmost liquid and simplest asset a company can have. It mainly refers to money in hand or in bank accounts that can be easily accessed.

Production Budget: A budget prepared in order to estimate the number of units that can be produced is called a production budget. It can be prepared either in a monthly format or a quarterly format.

Cash Budget: A budget prepared in order to estimate the inflows and outflows of cash in a particular period of time for a business organization is called a cash budget.

In order to sell the number of units 560,000, P company will have to manufacture budgeted number of units. Calculation of the budgeted number of units required to be produced are given below:

BudgetedProductionUnits=BudgetedSalesUnits+EndingInventoryofFGBeginningInventoryofFG=560,000units+66,000units96,000units=530,000units\begin{array}{c}\\{\rm{Budgeted Production Units}} = {\rm{Budgeted Sales Units}} + {\rm{Ending Inventory of FG}} - \\\\{\rm{Beginning Inventory of FG}}\\\\ = 560,000{\rm{ units}} + 66,000{\rm{ units}} - 96,000{\rm{ units}}\\\\ = 530,000{\rm{ units}}\\\end{array}

(2)

Computation of direct labor cost for the month of March is given below:

Totaldirecthourstobepaid=Guaranteedhoursforpayment×Laborrateperhour=5,480hours×$10=$54,800\begin{array}{c}\\{\rm{Total direct hours to be paid}} = {\rm{Guaranteed hours for payment }} \times \\\\{\rm{Labor rate per hour}}\\\\ = 5,480{\rm{ hours}} \times \$ 10\\\\ = \$ 54,800\\\end{array}

Working Note:

Computation of direct labor hours required for the month of March:

Totaldirecthoursrequired=BudgetedProductionunits×DirectLabourhoursrequiredperunit=7,100units×0.75=5,325hours\begin{array}{c}\\{\rm{Total direct hours required}} = {\rm{Budgeted Production units }} \times \\\\{\rm{Direct Labour hours required per unit}}\\\\ = 7,100{\rm{ units}} \times {\rm{0}}{\rm{.75}}\\\\ = 5,325{\rm{ hours}}\\\end{array}

Computation of direct labor cost for the month of March is given below:

Totaldirecthourstobepaid=Guaranteedhoursforpayment×Laborrateperhour=5,480hours×$10=$54,800\begin{array}{c}\\{\rm{Total direct hours to be paid}} = {\rm{Guaranteed hours for payment }} \times \\\\{\rm{Labor rate per hour}}\\\\ = 5,480{\rm{ hours}} \times \$ 10\\\\ = \$ 54,800\\\end{array}

Working Note:

Computation of direct labor hours required for the month of March:

Totaldirecthoursrequired=BudgetedProductionunits×DirectLabourhoursrequiredperunit=6,900units×0.75=5,175hours\begin{array}{c}\\{\rm{Total direct hours required}} = {\rm{Budgeted Production units }} \times \\\\{\rm{Direct Labour hours required per unit}}\\\\ = 6,900{\rm{ units}} \times {\rm{0}}{\rm{.75}}\\\\ = 5,175{\rm{ hours}}\\\end{array}

Computation of the total collective direct labor cost for a duration of two months is given below:

Collectivelaborhourstobepaid=LaborhourstobepaidinMarch+LaborhourstobepaidinApril=$54,800+$54,800=$109,600\begin{array}{c}\\{\rm{Collective labor hours to be paid}} = {\rm{Labor hours to be paid in March}} + \\\\{\rm{Labor hours to be paid in April}}\\\\ = \$ 54,800 + \$ 54,800\\\\ = \$ 109,600\\\end{array}

(3)

Computation of the amount that the company needs to be borrowed in order to maintain the closing cash balance that L company desires for the month of May is given below:

Amountneededtobeborrowed=DeriredcashbalanceAvailablecashbalance=$72,000$56,000=$16,000\begin{array}{c}\\{\rm{Amount needed to be borrowed}} = {\rm{Derired cash balance}} - {\rm{Available cash balance}}\\\\ = \$ 72,000 - \$ 56,000\\\\ = \$ 16,000\\\end{array}

Working Notes:

Computation of the available cash balance:

AvailableCashbalance=OpeningCashBalance+CashReceivedCashDisbursed=$51,000+$135,000$130,000=$56,000\begin{array}{c}\\{\rm{Available Cash balance}} = {\rm{Opening Cash Balance}} + {\rm{Cash Received}} - \\\\{\rm{Cash Disbursed}}\\\\ = \$ 51,000 + \$ 135,000 - \$ 130,000\\\\ = \$ 56,000\\\end{array}

Ans:

The budgeted number of units that P Company will have to produce in order to sell budgeted units 560,000, is 530,000 units.

Part 2

The collective labor hours for both March and April is $109,600.

Part 3

The amount of cash that is required to be borrowed by L company for the month of May is $16,000.